Obtain all the basic solutions of the following linear equations:
2x1 + x2 + x3 + x4 = 6
3x1 + 2x2 + x3 + 2x2 = 8
1
Expert's answer
2020-06-09T17:22:06-0400

"Since \ we \ have \\
2x_1 + x_2 + x_3 + x_4 = 6\\

3x_1 + 2x_2 + x_3 + 2x_4 = 8\\
This \ system \ can \ written\ in \ the \ form\\
\begin{bmatrix}
 2 & 1&1&1\\
 3&2&1&2
\end{bmatrix}
\begin{bmatrix}
 x_1\\x_2\\x_3\\x_4
\end{bmatrix}=\begin{bmatrix}
 6\\
 8
\end{bmatrix}\\
\text{Using Gauss elimination, we get }\\
\begin{bmatrix}
 \begin{array}{cccc|c}
 2 & 1 &1&1&6 \\ 
 3& 2 & 1&2&8 \\
 \end{array}
\end{bmatrix}
\xrightarrow {R_2-R_1\rightarrow R_2}\\ \\
\\
\begin{bmatrix}
 \begin{array}{cccc|c}
 2 & 1 &1&1&6 \\ 
 0& -1 &0&1&2 \\
 \end{array}
\end{bmatrix}\xrightarrow {-R_2\rightarrow R_2}\\ \\
\\
\begin{bmatrix}
 \begin{array}{cccc|c}
 3 & 1 &1&1&6 \\ 
 0& 1 &0&-1&-2 \\
 \end{array}
\end{bmatrix}\\
\text{So, we get the rank of matrix is 2}\\ \text{ and the number of unknowns is 4}\\
\text {this implies , we have 2 free unknown}\\
\text{So, let}\\
x_3=t , x_4=s.\\
\text{Now, we have}\\
x_2-x_4=4\Rightarrow x_2=4-s\\
2x_1 + x_2 + x_3 + x_4 = 6\\
\Rightarrow 2x_1 =- x_2 - t -s+ 6\\
2x_1 =-4+s - t -s+ 6\\
2x_1 =2-t\\"